Find the product or the quotient
37. 6 4/15 divided by (-2/3)
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

(6 4/15)/(-2/3) NOW CONVERT THE MIXED FRACTION 6 4/15 TO AN INPROPER FRACTION -
(6*15+4)/15=(90+4)/15=94/15.
NOW INVERT THE DENOMINATOR & MULTIPLY THESE FRACTIONS:
94/15*(-3/2)=-282/30=-47/5